OpenStudy (anikate):

if f(x)=(x^2) -1, and f(2a)=35, then what could be the value of a?

OpenStudy (anonymous):

Can you post a picture of this or dont copy and paste stuff since it doesn't make sense.

OpenStudy (anonymous):

@Anikate

OpenStudy (anikate):

@Austin1617 its #4

OpenStudy (anikate):

the answer is not B btw

OpenStudy (anonymous):

f(x) = 6^2 -1 f(x)=36-1 f(x) =35

OpenStudy (anonymous):

So E @Anikate

OpenStudy (anikate):

why??

OpenStudy (anikate):

oh ok wait

OpenStudy (anikate):

where did u get 6 from??

OpenStudy (mathmale):

f(x)=(x^2) -1 is the given function. Substitute (2a) for x, on both sides. Please show your work.

OpenStudy (anonymous):

I did simple math and 6 power of the 2 equals 36 then minus 1

OpenStudy (anikate):

sorry, still trying to understand this

OpenStudy (anikate):

sustitute 2a on both sides and what?

OpenStudy (mathmale):

f(x)=(x^2) -1 is the given function. Substitute (2a) for x, on both sides. Please show your work. Simplify your results as much as you can.

OpenStudy (mathmale):

I'm asking you to replace each "x" with "(2a)"

OpenStudy (mathmale):

f(x)=(x^2) -1 is the given function. Eliminating the "x" results in f( ) = ( )^2 - 1. Substitute (2a) into the each set of parentheses on both sides. Please show your work.

OpenStudy (mathmale):

I need your pretty much constant attention if you want to solve this problem.

OpenStudy (anikate):

ok yea i see

OpenStudy (mathmale):

f(x)=(x^2) -1 is the given function. Eliminating the "x" results in f( ) = ( )^2 - 1. Substitute (2a) into the each set of parentheses on both sides.

OpenStudy (anikate):

ok done

OpenStudy (anikate):

now?

OpenStudy (mathmale):

Type out your result, and then we'll discuss what to do next.

OpenStudy (anikate):

f(2a)=(2a)^2 -1

OpenStudy (mathmale):

Good. Evaluate (2a)^2 now, please. Then, subtract 1 from your result.

OpenStudy (anikate):

What do you mean by evaluate? what do i plug i for a? @mathmale

OpenStudy (mathmale):

Evaluate = find the value of ... This is algebra, so you need to become used to working with unknown quantities represented by letters, such as " a ". You don't plug in some value for " a ". .... Rather, you FIND the value of a. f(2a) = (2a)^2 - 1 = 35 (as given in the original problem statement). Please evaluate (find the equivalent expression of ) (2a)^2.

OpenStudy (mathmale):

f(2a) = [ ? ] - 1 =35

OpenStudy (anikate):

ooh ok so then a=(plus or minus) 3

OpenStudy (anikate):

a= 3 root 2

OpenStudy (mathmale):

I'd prefer to see all of your work, not just what you think the answer is. f(2a) = 4a^2 - 1 = 35. What did you do next?

OpenStudy (anikate):

add 1 to both sides, then divide by 4, so then u would get a^2 = 9, then a=3

OpenStudy (mathmale):

Right. Actually, a could be either 1) a=3 or 2) a=-3; both will satisfy the original equation.

OpenStudy (anikate):

-3 is the answer? but I was told that is wrong. its ok, I have to go to sleep now anyways, i'll double check with my instructor tomorrow

OpenStudy (mathmale):

You yourself could substitute a=-3 into the original equation. Then (2a)^2 = (2)^2 * (-3)^2 = +4(+9).